OpenStudy (theviper):

\(2\) cubes each of volume \(64cm^3\) are joined end to end. Find the surface area of the resulting shape.

OpenStudy (theviper):

i think this shape must be cuboid :/

OpenStudy (experimentx):

yeah .. it's a cuboid. first find the length of side of it.

OpenStudy (theviper):

Volume = \(a^3\) = \(64cm^3\) \(a\) = \(\sqrt{64cm}\) = \(4cm\)

OpenStudy (experimentx):

yep .. now let's visualize the picture of two cuboid joined together.

OpenStudy (anonymous):



OpenStudy (theviper):

so h=4cm, b=4cm but l=4+4=8cm ?

OpenStudy (anonymous):

Area = 4(s*2s)+2*s^2 where\[s=\sqrt[3]{64}=4 \]

OpenStudy (experimentx):

you know how to find surface area of one cuboid right??

OpenStudy (theviper):

yes :) Thanx everyone :)

OpenStudy (theviper):

You guys are so amazing ;)

OpenStudy (anonymous):

@experimentX Thank you for the medal.

OpenStudy (experimentx):

thank you for your effort too ...